### v1

@@ Abstract
Recent advancements in the Neural Radiance Field (NeRF) have bolstered its capabilities for novel view synthesis, yet its reliance on dense multi-view training images poses a practical challenge. Addressing this, we propose HourglassNeRF, an effective regularization-based approach with a novel hourglass casting strategy. Our proposed hourglass is conceptualized as a bundle of additional rays within the area between the original input ray and its corresponding reflection ray, by featurizing the conical frustum via Integrated Positional Encoding (IPE). This design expands the coverage of unseen views and enables an adaptive high-frequency regularization based on target pixel photo-consistency. Furthermore, we propose luminance consistency regularization based on the Lambertian assumption, which is known to be effective for training a set of augmented rays under the few-shot setting. Leveraging the inherent property of a Lambertian surface, which retains consistent luminance irrespective of the viewing angle, we assume our proposed hourglass as a collection of flipped diffuse reflection rays and enhance the luminance consistency between the original input ray and its corresponding hourglass, resulting in more physically grounded training framework and performance improvement. Our HourglassNeRF outperforms its baseline and achieves competitive results on multiple benchmarks with sharply rendered fine details. The code will be available.
Keywords:
Neural Radiance Field Few-shot Neural Rendering Ray Augmentation Hourglass Casting
@@ 1 Introduction
In recent years, the Neural Radiance Field (NeRF) [24] has emerged as a dominant paradigm in the domain of novel view synthesis, owing to its outstanding capability to produce high-quality rendered images and its inherently simplistic architectural design. Numerous advancements have been introduced to the NeRF [20, 3, 21, 23, 12, 43, 19, 18], propelling its performance to greater heights. However, its dependency on a dense set of multi-view training images still exists as a critical challenge for practical application.
@@
In the realm of few-shot novel view synthesis, two primary approaches have emerged as mainstream methodologies: pre-training and regularization. The pre-training methods [4, 5, 41, 22, 10, 17, 31, 38, 13, 46] necessitate extensive datasets comprising diverse scenes captured from multiple viewpoints, facilitating the infusion of prior knowledge for 3D geometry during the pre-training phase. However, it is highly expensive to collect large-scale datasets necessary for pre-training.
Mip-NeRF [1] RegNeRF [25] FlipNeRF [33] FreeNeRF [45] HourglassNeRF Ground Truth
Figure 1: Comparison with other baselines. Our HourglassNeRF achieves superior quality of fine details and textures to other state-of-the-art methods by applying the hourglass casting strategy equipped with an adaptive high-frequency regularization. The last two rows of the figure indicate the types of ray augmentation schemes and frequency regularization methods, respectively.
@@
In contrast, the regularization methods [25, 34, 14, 9, 32, 6, 16, 33, 45, 39, 35] are optimized to individual scenes, harnessing supplementary training assets, such as rays generated from unseen views [33, 25, 14, 16], pseudo-depth maps [6, 32, 39], readily available off-the-shelf models [9, 25], and so on. Moreover, these methods may directly introduce regularization techniques to the internal components of a NeRF framework, e.g. the density outputs of ray samples [34], the high-frequency segments stemming from positional encoding [45, 35], and so on. Although these methods have achieved promising results, they often rely on additional training resources, which may not be consistently accessible. Additionally, the existing frequency regularization technique requires the range of visible spectrum and the duration of masking phase, both of which should be set manually.
@@
Inspired by prior works [33, 45], we propose HourglassNeRF, which utilizes a novel hourglass casting strategy as an effective ray augmentation scheme. We cast an hourglass as a bundle of multiple additional rays, by featurizing the 3D Gaussian space of conical frustum using the Integrated Positional Encoding (IPE) [1]. This allows us to cover a broader area of unseen views as an additional training resource, further enhancing the efficiency of augmented rays. Furthermore, by applying IPE to the hourglass, we can adaptively apply high-frequency regularization based on the photo-consistency of the pixels where the original input ray and the newly generated hourglass are cast. It allows us to eliminate the hand-crafted aspects of the existing frequency regularization scheme and capture fine details better while preventing from overfitting to the high-frequency.
Our proposed HourglassNeRF is built upon FlipNeRF [33], which trains a set of augmented rays with existing ground-truth (GT) pixel values by taking advantage of Lambertian assumption. However, despite the Lambertian surface assumption, FlipNeRF utilized a set of flipped specular reflection rays, which goes against the physical characteristics of the training framework. Instead, we assume our proposed hourglass as a featurized bundle of flipped diffuse reflection rays originating from Lambertian surfaces, enabling a more physically reasonable training framework and leading to performance improvement. Furthermore, building upon the characteristic that a Lambertian surface maintains consistent luminance regardless of the viewing angle, we introduce luminance consistency regularization with an auxiliary luminance estimation task to enhance the consistency of estimated luminances between the original ray and its corresponding hourglass, which further improves the rendering quality as shown in Fig. 1. Our main contributions are summarized as follows:
@@
  • We propose a novel ray augmentation strategy, i.e. casting an hourglass as a bundle of additional rays, which is effective for training NeRF with only a set of sparse inputs.
@@
  • By marrying the IPE with our proposed hourglass, we can adaptively regularize the high-frequency of additional input samples based on the photo-consistency, which leads to performance improvement under the few-shot setting.
  • Furthermore, inspired by the fact that the Lambertian surface has the consistent luminance irrespective of viewing angle, we assume our proposed hourglass as multiple flipped diffuse reflection rays cast back from the Lambertian surface, and propose luminance consistency regularization for training our HourglassNeRF. By enhancing the consistency of estimated relative luminances between the original ray and the corresponding hourglass, we can further improve the integrity of our training framework and the rendering quality.
@@
  • Our HourglassNeRF outperforms its baseline, delivering competitive results across different datasets and scenarios with sharp fine details.
@@ 2.1 Neural Radiance Fields
The advent of NeRF [24] has facilitated a paradigm shift in novel view synthesis, achieving remarkable rendering quality. The NeRF employs a Multilayer Perceptron (MLP) to represent a scene, correlating spatial coordinates and viewing directions with corresponding color and volumetric density attributes, and then generates a novel view by volume rendering techniques. While NeRF’s versatility has been demonstrated across a gamut of applications—from the 3D object generation [20, 26, 21, 23] to capturing human performances [44, 48, 36, 12, 43], and even extending to dynamic videos [7, 19, 18, 37, 29]—its efficacy is notably hindered by a fundamental constraint: the reliance on a dense set of training images. This requirement poses significant challenges, particularly in scenarios restricted to a scant number of views, thus hampering its practical application. In addressing this critical bottleneck, our work focuses on improving NeRF’s performance under the constraints of limited input.
@@ 2.2 Few-shot Neural Rendering
In the emerging field of few-shot novel view synthesis, two principal strategies are prevalent: pre-training and regularization. The pre-training methods [4, 5, 41, 22, 10, 17, 31, 38, 13, 46] rely on large-scale datasets of multi-view scenes to infuse a NeRF model with a 3D geometry prior, often followed by fine-tuning on specific target scenes. In contrast, the regularization methods [25, 34, 14, 9, 32, 6, 16, 33, 45, 39, 35] are optimized to individual scenes by leveraging auxiliary training aids, e.g. pseudo-depth maps [6, 32, 39], augmented training rays [33, 25, 14, 16], semantic consistencies [9], and so on for additional guidance.
A notable technique within this approach, FlipNeRF [33], assumes the objects’ surface as Lambertian, which is effective in training extra data for few-shot NeRF, and exploits flipped specular reflection rays as additional training data, yielding promising outcomes. However, since Lambertian surfaces primarily exhibit diffuse reflection, using specular reflections goes against the physical traits expected under Lambertian assumptions, posing a challenge to the training framework’s integrity.
@@
In this work, we propose a novel hourglass casting strategy, which utilizes a featurized bundle of flipped diffuse reflection rays that conform to the physical expectations of Lambertian surfaces within our training framework. Furthermore, by applying IPE to featurize the hourglass, we can regulate the high-frequency of the additional hourglass adaptively based on the target pixel photo-consistency between the original input ray and the corresponding hourglass, leading to the improved rendering quality with clear fine details.
@@ 2.3 Ray Parameterization of NeRF
Since the original NeRF casts an infinitesimally narrow ray with point sampling, there has been a line of research exploring the effective ray shapes for higher-quality rendering, i.e. method to parameterize the ray [40, 1, 2, 8]. Among them, mip-NeRF [1] proposed a cone tracing strategy, in which each conical frustum area of the samples is featurized by the IPE, to tackle the aliasing problem. Mip-NeRF 360 [2] utilized contracted scene representation, which is an extended version of the mip-NeRF’s parameterization to unbounded scenes. Recently, Exact-NeRF [8] introduced a pyramidal parameterization method by using an analytically precise encoding technique instead of the IPE’s 3D Gaussian approximation, leading to performance improvement for unbounded scenes. In this paper, we propose a novel hourglass casting strategy as an effective ray augmentation scheme for few-shot NeRF, which we conceptualize as a featurized bundle of multiple rays, enabling an adaptive frequency regularization by combining with IPE. To the best of our knowledge, our HourglassNeRF is the first attempt to tackle the few-shot novel view synthesis task through the lens of ray parameterization.
@@ 3 Methods
Figure 2: Comparative overview of ray augmentation techniques. In contrast to other augmentation strategies ((b), (c)), where each additional ray corresponds to a single unseen view, our hourglass (d) encompasses a wider area of unseen views, thereby boosting the augmentation’s efficacy. Technically, the rays are spanned over the object surface as used in the typical NeRF’s volume rendering technique, but we simplify the illustration for easy comparison between the methods. Kindly refer to Fig. 3 for a more detailed description. The blue and red rays indicate the original and augmented ray, respectively.
In this work, we propose HourglassNeRF, an effective regularization method for few-shot novel view synthesis with an hourglass casting strategy. Our HourglassNeRF is built upon FlipNeRF, which utilizes flipped reflection rays with Lambertian assumption (Sec. 3.1). We generate a set of hourglasses whose conical frustums are featurized by IPE and cast them toward the identical target pixels as additional training rays, adaptively regulating their high-frequency components (Sec. 3.2). Furthermore, taking advantage of the characteristics of the Lambertian assumption, we propose the luminance consistency regularization to enhance the consistency of the estimated relative luminances between the original ray and its corresponding hourglass, leading to further performance improvement (Sec. 3.3). A brief comparison with different ray augmentation schemes is shown in Fig. 2.
@@ 3.1 Preliminaries
Mip-NeRF.
The mip-NeRF [1], which casts cones instead of rays to alleviate aliasing, represents a 3D scene by mapping 3D coordinates $\mathbf{x}=(x,y,z)$ and viewing directions $(\theta,\phi)$ to the colors $\mathbf{c}=(r,g,b)$ and volumetric densities $\tau$:
@@
FlipNeRF.
Implemented upon mip-NeRF, the FlipNeRF [33] utilizes a set of flipped reflection rays $\mathbf{r}^{\prime}$ as additional training rays, which are derived from the original ray directions $\mathbf{\bar{d}}$ and the estimated normal vectors $\mathbf{\hat{n}}=\sum_{i=1}^{N}w_{i}\mathbf{n}_{i}$ as follows:
@@
where $\mathbf{d}^{\prime}=2(\mathbf{\bar{d}}\cdot\hat{\mathbf{n}})\hat{\mathbf{n}}-\mathbf{\bar{d}}$ and $\mathbf{o}^{\prime}=\mathbf{p}_{s}-t_{s}\mathbf{d}^{\prime}$ indicate the flipped reflection direction and the newly set camera origin derived from the object’s estimated surface $\mathbf{p}_{s}=\mathbf{o}+t_{s}\mathbf{\bar{d}}$, respectively. Note that $t_{s}$ is the distance to the estimated object surface. Taking advantage of Lambertian surface assumption, which is useful for a few-shot training framework utilizing the GT pixel values of original rays to provide relevant supervisory signals for additional rays, the flipped reflection rays can be optimized effectively with filtering the ineffective rays whose angle with the corresponding original rays are over the masking threshold $\psi=90^{\circ}$. However, despite the Lambertian assumption, the FlipNeRF makes use of specular reflection rays, which are mainly originated from the non-Lambertian surface, making its framework conflict with physical characteristics.
@@
Our HourglassNeRF is built upon FlipNeRF while achieving better rendering quality with more fine details compared to FlipNeRF thanks to our proposed hourglass casting. Furthermore, we conceptualize the hourglass as a bundle of flipped diffuse reflection rays in line with the Lambertian assumption, leading to a more physically reasonable training framework.
3.2 Hourglass Casting
[…]
@@ 3.2 Hourglass Casting
Figure 3: Hourglass generation process. (a) First, to derive $\tilde{\sigma}^{2}_{\rho}$, we reparameterize the original metric distance $t$ as $\tilde{t}$. We shift $t_{1}$, i.e. starting point of a ray along a z-axis, to the estimated object surface $\mathbf{p}_{s}$, so that our proposed hourglass is symmetrically constructed around $\mathbf{p}_{s}$, i.e. $\tilde{t}_{s-i}=\tilde{t}_{s+i}$ and it leads to $\boldsymbol{\tilde{\Sigma}}_{s-i}=\boldsymbol{\tilde{\Sigma}}_{s+i}$. Note that $\tilde{t}$ is used only for $\tilde{\sigma}^{2}_{\rho}$. (b) Using the trigonometric function, we compute the base radius of hourglass $\tilde{\rho}$. As a result, our proposed hourglass $\mathbf{\tilde{r}}$ is featurized to cover the unseen view area between the original ray and its reflection ray. The blue dotted ray denotes the reflection ray of $\mathbf{r}$.
@@
We propose an hourglass-shaped ray as an effective additional training ray for few-shot NeRF. As shown in Fig. 2, compared to the existing ray augmentation schemes where a resulting augmented ray corresponds to an unseen view, our proposed hourglass covers the area of continuous unseen views by IPE presented in Eq. 2, providng more efficient extra training resources.
@@
First, as demonstrated in Fig. 3, we reparameterize the metric distance $t\in[t_{near},$ $t_{far}]$ as $\tilde{t}$ to derive the variance $\tilde{\sigma}^{2}_{\rho}$, which is perpendicular to the hourglass. We shift $t$ so that the value of starting point $t_{1}$ is located on the distance to the estimated object surface $t_{s}$, resulting in $\tilde{t}_{s}=t_{1}$. And then we set $[t_{1},t_{s-1}]$ values to be symmetric with respect to $t_{s}$.
@@
Next, as shown in Fig. 3, we derive a base radius of the hourglass $\tilde{\rho}$ from the angle $\tilde{\theta}$ between $\mathbf{r}$ and $\mathbf{\hat{n}}$ using the trigonometric function as follows:
@@
where $\tilde{\delta}=1-\tilde{t}_{s}$ so that $\tilde{\rho}$ is obtained from the sample located on $\tilde{t}=1$ following [1]. However, directly employing the obtained $\tilde{\rho}$ in IPE results in significantly large $\tilde{\sigma}^{2}_{\rho}$, leading to over-regularization of high-frequency components for the samples along an hourglass. Thus, we adjust the scale of $\tilde{\rho}$ to $[0,1]$ to contract $\tilde{\rho}$ with the large $\tilde{\theta}$ value into a proper range, while leaving the one with small $\tilde{\theta}$ affected little as follows:
@@
This ensures high-frequency regularization is appropriately applied.
And then, $\tilde{\sigma}^{2}_{\rho}$ is derived from $\tilde{t}$ and $\tilde{\rho}$ to featurize the conical frustums of hourglass as multivariate Gaussian by simply replacing the original metric distance $t$, which is used in mip-NeRF, with $\tilde{t}$ as follows:
@@
where $\tilde{t}_{\delta}$ and $\tilde{t}_{\mu}$ denote a half-width and mid-point of adjacent $\tilde{t}$ values. Note that we use the same $\mu_{t}$ and $\sigma^{2}_{t}$ for the mean and variance along the hourglass as mip-NeRF.
Finally, we generate an hourglass $\mathbf{\tilde{r}}(t)=\mathbf{\tilde{o}}+t\mathbf{\tilde{d}}$, where $\mathbf{\tilde{d}}=-\mathbf{\hat{n}}$ and $\mathbf{\tilde{o}}=\mathbf{p}_{s}-t_{s}\mathbf{\tilde{d}}$, so that the hourglass is cast from the newly set camera origin $\mathbf{\tilde{o}}$, which has the same distance from $\mathbf{p}_{s}$ as the original ray, covering the unseen view area between the original ray and the corresponding reflection ray around the axis of $\mathbf{\hat{n}}$. Implemented upon FlipNeRF, our HourglassNeRF is optimized with the same training losses while using our proposed hourglasses as an additional training batch instead of flipped reflection rays.
Since the target pixel photo-consistency between $\mathbf{r}$ and $\mathbf{\tilde{r}}$ varies depending on the angle $\tilde{\theta}$ between them, which is used to derive the base radius of hourglass $\tilde{\rho}$ and the x/y-axis variance $\tilde{\sigma}^{2}_{\rho}$, the high-frequency spectrum of samples along $\mathbf{\tilde{r}}$ are regulated adaptively based on the photo-consistency, i.e. the less photo-consistent the target pixel is, the more regulated the samples’ high-frequency components are. It eliminates the need for manually designing a frequency spectrum to regularize during the training phase and effectively adjusts the amount of high-frequency detail retained in additional training samples for learning fine details.

### v2

@@ Abstract
footnotetext: §Corresponding author.
Recent advancements in the Neural Radiance Field (NeRF) have enhanced its capabilities for novel view synthesis, yet its reliance on dense multi-view training images poses a practical challenge, often leading to artifacts and a lack of fine object details. Addressing this, we propose ARC-NeRF, an effective regularization-based approach with a novel Area Ray Casting strategy. While the previous ray augmentation methods are limited to covering only a single unseen view per extra ray, our proposed Area Ray covers a broader range of unseen views with just a single ray and enables an adaptive high-frequency regularization based on target pixel photo-consistency. Moreover, we propose luminance consistency regularization, which enhances the consistency of relative luminance between the original and Area Ray, leading to more accurate object textures. The relative luminance, as a free lunch extra data easily derived from RGB images, can be effectively utilized in few-shot scenarios where available training data is limited. Our ARC-NeRF outperforms its baseline and achieves competitive results on multiple benchmarks with sharply rendered fine details.
Mip-NeRF [1] RegNeRF [26] FlipNeRF [34] FreeNeRF [50] ARC-NeRF Ground Truth
Figure 1: Comparison with other baselines. Our ARC-NeRF achieves superior quality of fine details and textures to other state-of-the-art methods by using our Area Rays equipped with an adaptive high-frequency regularization. The last two rows of the figure indicate the types of ray augmentation schemes and frequency regularization methods, respectively.
@@ 1 Introduction
In recent years, the Neural Radiance Field (NeRF) [25] has emerged as a dominant paradigm in the domain of novel view synthesis, owing to its outstanding capability to produce high-quality rendered images and its inherently simplistic architectural design. Numerous advancements have been introduced to the NeRF [21, 3, 22, 24, 12, 46, 20, 19], taking its performance to the next level. However, its dependency on a dense set of multi-view training images still exists as a critical challenge for practical application, resulting in floating artifacts and missing fine details in object textures.
@@
Two primary approaches have emerged as mainstream methodologies for the few-shot novel view synthesis: pre-training and regularization. The pre-training methods [4, 5, 44, 23, 10, 18, 32, 40, 13, 51] necessitate extensive datasets comprising diverse scenes captured from multiple viewpoints, facilitating the infusion of prior knowledge for 3D geometry during the pre-training phase. However, it is highly expensive to collect large-scale datasets necessary for pre-training.
@@
In contrast, the regularization methods [26, 35, 15, 9, 33, 6, 17, 34, 50, 41, 37, 43, 36, 47, 48] are optimized per scene, using extra training assets, such as rays generated from unseen views [34, 26, 15, 17], pseudo-depth maps [6, 33, 41, 43], readily available off-the-shelf models [9, 26, 47], and so on. Moreover, these methods may directly apply regularization techniques to the internal components of a NeRF framework, e.g. the density outputs of ray samples [35] and the high-frequency segments stemming from positional encoding [50, 37]. Although these methods have achieved promising results, they often rely on extra training resources, which may not be consistently accessible. Also, the existing frequency regularization technique requires the range of masked frequency and the duration of masking phase, both of which should be set manually.
@@
To address the problem, we propose ARC-NeRF, which utilizes a novel Area Ray casting strategy as an effective ray augmentation scheme. We cast an Area Ray as a bundle of multiple additional rays, by featurizing the 3D Gaussian space of conical frustum using the Integrated Positional Encoding (IPE) [1]. This enables us to cover a broader area of unseen views as an additional training resource, enhancing the efficiency of augmented rays, whereas previous ray augmentation methods are limited to covering only a single unseen view per additional ray.
Furthermore, by applying IPE to the Area Ray, we can adaptively apply high-frequency regularization based on the photo-consistency of the pixels where the original input ray and the newly generated Area Ray are cast. Since the photo-consistency varies depending on the angle between the original and Area Ray, which determines the radius and variance of Area Ray, the high-frequency components of Area Ray samples are adaptively regulated based on the photo-consistency. It allows us to eliminate the hand-crafted aspects of the existing frequency regularization scheme and capture fine details better while preventing from overfitting on the high-frequency.
Additionally, we propose luminance map, which serves as free lunch additional data that can be easily derived from RGB images, as an effective additional training resource in few-shot scenarios with limited training data, and introduce luminance consistency regularization with an auxiliary luminance estimation task. It enhances the consistency of estimated luminance between the original ray and its corresponding Area Ray, allowing for more accurate capture of delicate object textures with fine details as shown in Fig. 1. Our main contributions are as follows:
@@
  • We propose a novel ray augmentation strategy, i.e. casting an Area Ray as a bundle of additional rays to cover a broader range of unseen views, which is effective for training NeRF with only a set of sparse inputs.
@@
  • By marrying the IPE with our Area Ray, we can adaptively regularize the high-frequency of additional input samples based on the photo-consistency, leading to improved detail in rendering under the few-shot scenario.
  • We propose a luminance map, easily derived from RGB images as simple but effective free lunch data, and introduce luminance consistency regularization through an auxiliary luminance estimation task, further enhancing the rendering quality.
@@
  • Our ARC-NeRF outperforms its baseline, delivering competitive results across different datasets and scenarios with sharp fine details in object textures.
@@ 2.1 Neural Radiance Fields
The advent of NeRF [25] has facilitated a paradigm shift in novel view synthesis, achieving remarkable rendering quality. The NeRF employs a Multilayer Perceptron (MLP) to represent a scene, correlating 3D coordinates and viewing directions with corresponding color and volumetric density attributes, and then generates a novel view by volume rendering techniques. While NeRF’s versatility has been demonstrated across various applications, e.g. the 3D object generation [21, 27, 22, 24], capturing human performances [49, 53, 38, 12, 46], dynamic videos [7, 20, 19, 39, 30], and so on, its efficacy is notably hindered by a fundamental constraint: the reliance on a dense set of training images. This requirement poses significant challenges, particularly in scenarios restricted to a scant number of views, thus hampering its practical application. In addressing this critical bottleneck, our work focuses on improving NeRF’s performance under the constraints of limited input.
@@ 2.2 Few-shot Neural Rendering
In the emerging field of few-shot novel view synthesis, two principal strategies are prevalent: pre-training and regularization. The pre-training methods [4, 5, 44, 23, 10, 18, 32, 40, 13, 51] rely on large-scale datasets of multi-view scenes to infuse a NeRF model with a 3D geometry prior, often followed by fine-tuning on specific target scenes. In contrast, the regularization methods [26, 35, 15, 9, 33, 6, 17, 34, 50, 41, 37, 36, 47, 43, 48] are optimized to individual scenes by leveraging auxiliary training aids, e.g. pseudo-depth maps [6, 33, 41, 43], augmented training rays [34, 26, 15, 17], semantic consistencies [9], and so on for additional guidance.
Among them, the ray augmentation methods [34, 26, 15, 17] directly address the lack of training data by generating additional rays for training, effectively tackling data limitations. However, since the number of augmented rays directly increases the training cost, generating a large number of additional rays is practically challenging. Therefore, the existing ray augmentation methods have limitations in fully utilizing available unseen views. A notable technique within this approach, FlipNeRF [34] exploits flipped reflection rays as additional training data, yielding promising outcomes. However, since only a single reflection ray is used per original ray for an unseen view, it still does not leverage the potentially effective additional rays that exist between them from unseen views.
@@
In this work, we propose a novel Area Ray casting strategy that uses a featurized bundle of multiple rays to cover not just a single unseen view, but a broad range of unseen views. Furthermore, by applying IPE to featurize the Area Ray, we can regulate the high-frequency of the Area Ray adaptively based on the target pixel photo-consistency between the original input ray and the corresponding Area Ray, leading to the improved quality with clear fine details.
@@ 2.3 Ray Parameterization of NeRF
Since the original NeRF casts an infinitesimally narrow ray with point sampling, there has been a line of research exploring the effective ray shapes for higher-quality rendering, i.e. method to parameterize the ray [42, 1, 2, 8]. Among them, mip-NeRF [1] proposed a cone tracing strategy, in which each conical frustum area of the samples is featurized by the IPE, to tackle the aliasing problem. Mip-NeRF 360 [2] utilized contracted scene representation, which is an extended version of the mip-NeRF’s parameterization to unbounded scenes. Recently, Exact-NeRF [8] introduced a pyramidal parameterization method by using an analytically precise encoding technique instead of the IPE’s 3D Gaussian approximation, leading to performance improvement for unbounded scenes. In this paper, we propose a novel Area Ray casting strategy as an effective ray augmentation scheme for few-shot NeRF, which we conceptualize as a featurized bundle of multiple rays, covering wider unseen view areas and enabling an adaptive frequency regularization by combining with IPE. To the best of our knowledge, our ARC-NeRF is the first to tackle the few-shot novel view synthesis task through the lens of ray parameterization.
@@ 3 Methods
Figure 2: Comparative overview of ray augmentation techniques. In contrast to other augmentation strategies ((b), (c)), where each additional ray corresponds to a single unseen view, our Area Ray (d) encompasses a wider area of unseen views, thereby boosting the augmentation’s efficacy. Technically, the rays are spanned over the object surface as used in the typical NeRF’s volume rendering technique, but we simplify the illustration for easy comparison between the methods. Kindly refer to Fig. 3 for a more detailed description. The blue and red rays indicate the original and augmented ray, respectively.
In this work, we propose an effective regularization method for few-shot novel view synthesis with an Area Ray casting strategy. Our ARC-NeRF is built upon FlipNeRF, which utilizes flipped reflection rays as the augmented rays (Sec. 3.1). Instead of the flipped reflection rays, we generate a set of Area Rays whose conical frustums are featurized by IPE and cast them toward the identical target pixels as additional training rays, adaptively regulating their high-frequency components (Sec. 3.2). Furthermore, by taking advantage of the luminance map, i.e. free lunch data that can be easily derived from RGB images, we propose the luminance consistency regularization to enhance the consistency of the estimated relative luminances between the original ray and its corresponding Area Ray, leading to further performance improvement (Sec. 3.3). A brief comparison with different ray augmentation schemes is shown in Fig. 2.
@@ 3.1 Preliminaries
Mip-NeRF.   The mip-NeRF [1], which casts cones instead of rays to alleviate aliasing, represents a 3D scene by mapping 3D coordinates $\mathbf{x}=(x,y,z)$ and viewing directions $(\theta,\phi)$ to the colors $\mathbf{c}=(r,g,b)$ and volumetric densities $\tau$:
@@
FlipNeRF.   Implemented upon mip-NeRF, the FlipNeRF [34] utilizes a set of flipped reflection rays $\mathbf{r}^{\prime}$ as additional training rays, which are derived from the original ray directions $\mathbf{\bar{d}}$ and the estimated normal vectors $\mathbf{\hat{n}}=\sum_{i=1}^{N}w_{i}\mathbf{n}_{i}$ as follows:
@@
where $\mathbf{d}^{\prime}=2(\mathbf{\bar{d}}\cdot\hat{\mathbf{n}})\hat{\mathbf{n}}-\mathbf{\bar{d}}$ and $\mathbf{o}^{\prime}=\mathbf{p}_{s}-t_{s}\mathbf{d}^{\prime}$ indicate the flipped reflection direction and the newly set camera origin derived from the object’s estimated surface $\mathbf{p}_{s}=\mathbf{o}+t_{s}\mathbf{\bar{d}}$, respectively. Note that $t_{s}$ is the distance to the estimated object surface. By utilizing the GT pixel values of original rays to provide relevant supervisory signals for additional rays, the flipped reflection rays can be optimized effectively with filtering the ineffective rays whose angle with the corresponding original rays are over the masking threshold $\psi=90^{\circ}$. However, despite the countless potential candidates for effective additional rays in unseen view areas under the masking threshold, only the flipped reflection ray is utilized, leading to suboptimal training efficiency.
@@
Our ARC-NeRF is built upon FlipNeRF while achieving better rendering quality with more fine details compared to FlipNeRF thanks to our proposed Area Ray casting. Furthermore, we conceptualize the Area Ray as a bundle of multiple rays to cover a broad unseen view region, more optimally utilizing ray augmentation.
3.2 Area Ray Casting
(a) Metric distance reparameterization
(b) Area Ray featurization by IPE
@@ 3.2 Area Ray Casting
Figure 3: Area Ray generation process. (a) First, to derive $\tilde{\sigma}^{2}_{\rho}$, we reparameterize the original metric distance $t$ as $\tilde{t}$. We shift $t_{1}$, i.e. starting point of a ray along a z-axis, to the estimated object surface $\mathbf{p}_{s}$, so that our proposed Area Ray is symmetrically constructed around $\mathbf{p}_{s}$, i.e. $\tilde{t}_{s-i}=\tilde{t}_{s+i}$ and it leads to $\boldsymbol{\tilde{\Sigma}}_{s-i}=\boldsymbol{\tilde{\Sigma}}_{s+i}$. Note that $\tilde{t}$ is used only for $\tilde{\sigma}^{2}_{\rho}$. (b) Using the trigonometric function, we compute the base radius of Area Ray $\tilde{\rho}$. As a result, our proposed Area Ray $\mathbf{\tilde{r}}$ is featurized to cover the unseen view area between the original ray and its reflection ray. The blue dotted ray denotes the reflection ray of $\mathbf{r}$.
@@
We propose an Area Ray as an effective additional training ray for few-shot NeRF. As shown in Fig. 2, compared to the existing ray augmentation schemes where a resulting augmented ray corresponds to an unseen view, our proposed Area Ray covers the area of continuous unseen views by IPE presented in Eq. 2, providng more efficient extra training resources.
@@
First, we reparameterize the metric distance $t\in[t_{near},$ $t_{far}]$ as $\tilde{t}$ to derive the variance $\tilde{\sigma}^{2}_{\rho}$, which is perpendicular to the Area Ray (Fig. 3(a)). We shift $t$ so that the value of starting point $t_{1}$ is located on the distance to the estimated object surface $t_{s}$, resulting in $\tilde{t}_{s}=t_{1}$. And then we set $[t_{1},t_{s-1}]$ values to be symmetric with respect to $t_{s}$.
@@
Next, as shown in Fig. 3(b), we derive a base radius of the Area Ray $\tilde{\rho}$ from the angle $\tilde{\theta}$ between $\mathbf{r}$ and $\mathbf{\hat{n}}$ using the trigonometric function as follows:
@@
where $\tilde{\delta}=1-\tilde{t}_{s}$ so that $\tilde{\rho}$ is obtained from the sample located on $\tilde{t}=1$ following [1]. However, directly employing the obtained $\tilde{\rho}$ in IPE results in significantly large $\tilde{\sigma}^{2}_{\rho}$, leading to over-regularization of high-frequency components for the samples along an Area Ray. Thus, we adjust the scale of $\tilde{\rho}$ to $[0,1]$ to contract $\tilde{\rho}$ with the large $\tilde{\theta}$ value into a proper range, while leaving the one with small $\tilde{\theta}$ affected little as follows:
@@
And then, $\tilde{\sigma}^{2}_{\rho}$ is derived from $\tilde{t}$ and $\tilde{\rho}$ to featurize the conical frustums of Area Ray as multivariate Gaussian by simply replacing the original metric distance $t$, which is used in mip-NeRF, with $\tilde{t}$ as follows:
@@
where $\tilde{t}_{\delta}$ and $\tilde{t}_{\mu}$ denote a half-width and mid-point of adjacent $\tilde{t}$ values. Note that we use the same $\mu_{t}$ and $\sigma^{2}_{t}$ for the mean and variance along the Area Ray as mip-NeRF.
Finally, we generate an Area Ray $\mathbf{\tilde{r}}(t)=\mathbf{\tilde{o}}+t\mathbf{\tilde{d}}$, where $\mathbf{\tilde{d}}=-\mathbf{\hat{n}}$ and $\mathbf{\tilde{o}}=\mathbf{p}_{s}-t_{s}\mathbf{\tilde{d}}$, so that the Area Ray is cast from the newly set camera origin $\mathbf{\tilde{o}}$, which has the same distance from $\mathbf{p}_{s}$ as the original ray, covering the unseen view area between the original ray and the corresponding reflection ray around the axis of $\mathbf{\hat{n}}$. Implemented upon FlipNeRF, our ARC-NeRF is optimized with the same training losses while using our proposed Area Rays as an additional training batch instead of flipped reflection rays.
Since the target pixel photo-consistency between $\mathbf{r}$ and $\mathbf{\tilde{r}}$ varies depending on the angle $\tilde{\theta}$ between them, which is used to derive the base radius of Area Ray $\tilde{\rho}$ and the x/y-axis variance $\tilde{\sigma}^{2}_{\rho}$, the high-frequency spectrum of samples along $\mathbf{\tilde{r}}$ are regulated adaptively based on the photo-consistency, i.e. the less photo-consistent the target pixel is, the more regulated the samples’ high-frequency components are. It eliminates the need for manually designing a frequency spectrum to regularize during the training phase and effectively adjusts the amount of high-frequency detail retained in additional training samples for learning fine details.